OpenStudy (anonymous):

The monthly revenue of a certain company is given by R=800p-7p^2 where p is the price in dollars of the product the company manufactures. At what price will the revenue be $10000 if the price must be greater than $50?

OpenStudy (gw2011):

800p-7p^2=10,000 -7p^2+800p-10,000=0 (-7p+100)(p-100)=0 p-100=0 p=100 -7p+100=0 -7p=-100 p=-100/-7=14.29 Since the price must be greater than $50, the price then is $100

OpenStudy (anonymous):

thxs

OpenStudy (gw2011):

You're welcome